OpenStudy (anonymous):

how would you solve 4!

OpenStudy (saifoo.khan):

Calc.

OpenStudy (anonymous):

n! = n*(n-1)*(n-2)*...*2*1 so you just count down to 1 and multiply everything. 4! = 4*3*2*1 = 24

OpenStudy (gw2011):

4! = 4 factorial = (4)(3)(2)(1) = 24 Factorial means that you take that number and go down in decreasing order and then you multiply all those numbers.